OpenStudy (anonymous):

What is the exact value of 14 times z squared all over the square root of z to the ninth power.?

OpenStudy (anonymous):

\[\frac{ 14*z^{2} }{ \sqrt{z^{9}} }\]

OpenStudy (anonymous):

that it

OpenStudy (anonymous):

which one: 14 the square root of z all over z to the seventh power. 14 the square root of z all over z cubed. 14 the square root of z all over z squared. 14 over z squared.

OpenStudy (anonymous):

One moment

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\Large \frac{ 14 \times z^{2} }{ \sqrt{z^{9}}}\] \[\Large \frac{ 14 \times z^{2} }{ z^\frac{1}{9} }\] \[\Large 14 \times z^{2-\frac{1}{9}}\]

OpenStudy (anonymous):

so thats D

OpenStudy (anonymous):

Not exactly